The base of an isosceles triangle is 21 centimeters long. The altitude to the base is 9 centimeters long. What is the approximate measure of a base angle of the triangle
Answer by RAY100(1637)

A rough sketch helps to visualize.
.
the base of the isosceles triangle is 21, and the median from the opposite vertex is the height,,, which is 9
.
This causes a right triangle, legs 9 and 10.5 (21/2)
.
The base angle of the triangle has, tan x = opp/adj = h/(b/2) =9/10.5 = .8571
.
the angle x = tan^-1 (.8571) = 40.60 degrees
.
.
check....9/10.5 = tan 40.6,,,,.8571,,,,,ok